Find the center and radius of the circle with equation (x - 6)2 + (y + 5)2 = 9
Answer by Alan3354(69443)

Find the center and radius of the circle with equation (x - 6)2 + (y + 5)2 = 9
========================
A circle - (x-h)^2 + (y-k)^2 = r^2 has its center at (h,k) and radius r.